# Interstellar comet 3I/ATLAS: discovery and physical description

### v2

@@ Abstract
We describe the physical characteristics of interstellar comet 3I/ATLAS, discovered on 2025 July 1 by the Asteroid Terrestrial-impact Last Alert System. The comet has eccentricity, 6.08 and velocity at infinity, v∞ 57 km/s, indicating an interstellar origin. We obtained B,V, R, I, g, r, i, and z photometry with the Kottamia Astronomical Observatory 1.88-m, the Palomar 200-inch, and the Astrophysical Research Consortium 3.5-m on 2024 July 2, 3, and 6. We measured colour indices B–V=0.980.23, V–R=0.710.09, R–I=0.140.10, g–r=0.840.05 mag, r–i=0.160.03 mag, i–z=–0.020.07 mag, and g–i=1.000.05 mag and a spectral slope of 16.01.9 /100 nm. We calculate the dust cross-section within 10,000 km of the comet to be 184.64.6 km2, assuming an albedo of 0.10. 3I/ATLAS’s coma has FWHM2.2 arcsec and A(0∘)f=280.83.2 cm. We estimate that 3I/ATLAS’s µm-scale to mm-scale dust is ejected at 0.01-1 m/s, implying a dust production of 0.1 - 1.0 kg/s.
@@ 2 Observations
Initially designated as A11pl3Z, interstellar comet 3I/ATLAS was discovered at the ATLAS Rio Hurtado, Chile facility (Minor Planet Center (MPC) observing code W68). The initial discovery observations of 3I/ATLAS were made on 2025 July 1 05:15:11 in the ATLAS o-band filter as seen in Panels a-d of Figure 1 (Denneau et al., 2025). The comet was discovered at right ascension, declination of 18:07:27.68 and -18:41:40.2 and was detected in four o-band exposures taken between 2025 July 1 05:15:11and 2025 July 1 06:20:31. A list containing details of the ATLAS observations is presented in Table S1 of the Supplemental Material.
@@
Images from all three facilities were reduced using biases and flats taken for each of the filters. Image stacks of the NGPS and ARCTIC data are shown in the mosaic images in Figs. 2 and 3. The multi-band data were calibrated using the Pan-STARRS1 catalog (Chambers et al., 2016) and the colour transformations from (Tonry et al., 2012) to convert Pan-STARR1 photometry magnitudes into SDSS photometry magnitudes. The photometry of 3I/ATLAS was measured using a 4″ radius circular aperture (10,000 km from the comet’s nucleus at its 3.36 - 3.45 au distance from the Earth during our KAO 1.88-m, P200, and ARCTIC 3.5-m observaitons). We used a 6.5-7.8″ radius sky background annulus to measure our photomery. The seeing at the KAO 1.88-m, the P200, and the ARC 3.5-m measured between 1.3-1.6 arcsec in images containing the comet. A Table containing details of the KAO 1.88-m, P200, and ARC 3.5-m observations is presented in Table S1 of the Supplemental Material.
@@ 3.1 Orbit
The astrometry of 3I/ATLAS from ATLAS taken in the discovery images on 2025 July 1 was reported to the MPC. Astrometry of the comet in 319 observations taken from 97 observational facilities around the world between 2024 June 14.25 UTC and 2025 July 4 UTC was subsequently reported to the MPC 111https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=3I, accessed on 2025 July 4.. The orbital solution from JPL HORIZONS, using observations taken between June 14, 2025, and July 4, 2025, is shown in Table S2. The location and orbit of 3I/ATLAS at the time of its discovery on 2024 July 4 is shown on Figure 4 based on values from JPL Horizons222https://ssd.jpl.nasa.gov/horizons/app.html#/.
@@ 3.2 Visible photometry
We computed the brightness of 3I/ATLAS in the KAO 1.88-m images of B=18.90 0.21, V=17.92 0.01, R=17.22 0.05, and I=17.08 0.09, in the NGPS images of g=18.720.02, r=17.860.01, and i=17.720.02, and in the ARCTIC images of g=18.580.06, r=17.770.02, i=17.590.04, and z=17.610.05. We used the NGPS r-band stack to search for evidence of an extended coma in 3I/ATLAS. The surface brightness profile extracted from the r-band images of 3I/ATLAS (blue line in Panel D of Figure 2) shows evidence of being extended, having 2-10 times the flux compared to the surface brightness profile of a comparison star (orange line in the same panel) outside of 1.3 arcsec of both profiles. The FWHM of the coma in the NGPS r-band stack is 2.2 arcsec compared to the 1.6 arcsec FWHM of background stars in the sidereal-tracked images. 3I/ATLAS ’s colours are B–V=0.980.23, V–R=0.710.09, R–I=0.140.10, g–r=0.840.05 mag, r–i=0.160.03 mag, i–z=–0.020.07 mag, and g–i=1.000.05 mag corresponding to a spectral slope of 16.01.9 /100 nm., somewhat more red than interstellar comet 2I/Borisov which had g–r=0.630.05, r–i=0.200.02, g–i=0.830.05, and a spectral slope of 9.71.8 /100 nm. The colors of 3I/ATLAS derived from our observations are similar to those reported by other photometric and spectroscopic studies (e.g., Opitom et al., 2025). We plot the g–r and r–z colours of 3I/ATLAS in Figure 5.
@@
where is the half-width half maximum measured from our r-band stack of 1.1 arcsec, is the measured half-width half maximum of background stars of 0.8 arcsec, resulting in a of 0.7 arcsec, which translates to 1850 km at the 3.43 au distance of the comet. Plugging our calculated value of of 1850 km into Eq. 4 gives 1.1m s-1 suggesting that m-scale to mm-scale particles are ejected from the comet at 0.01-1 m/s. The dust mass loss, can be roughly estimated from , A(0∘)f, the dust albedo , the radius of the dust particles, , and the dust density, , with the following relation derived from (Fink & Rubin, 2012, see also Bernardinelli et al. (2021)):
@@ 4 Discussion and conclusion
Interstellar comet 3I/ATLAS exhibits an active appearance and colours similar to some solar system comets, as seen in Figures 2 and 5. However, it seems to be significantly more red in its color compared to the other interstellar comet 2I (g–r=0.630.05, r–i=0.200.02, and g–i=0.830.05), however, is less red compared to 1I/’Oumuamua (g–r=0.840.05, r–i=0.360.1, and g–i=1.200.11). However, this latter comparison may be inaccurate due to the fact that 3I/ATLAS is an active body, and 1I/’Oumuamua was only very weakly active as evidenced by the visible lack of a coma (Micheli et al., 2018). A complete table of the physical properties of the different interstellar objects is included in Table S3. Interstellar comet 3I/ATLAS appears to be more red than the majority of comets as seen in Fig. 4. The increased redness of 3I compared to 2I and other active comets may be due to the nature of its activity if it lacks the emission of gas molecules present in the g-band wavelength such as CN, C2, and C2 (Farnham et al., 2000). With our small sample size of three objects, it appears that the interstellar objects span a wide range in color. Additionally, 3I shows evidence of having a similar dust mass loss compared to 2I, but has a much higher dust mass loss rate than 1I, which was observed to show a negligible amount of dust mass loss.
@@ Acknowledgements
et al., 1998). The seeing in r-band was 1.6 arcsec measured by determining the full-width at half-maximum of background stars during the observations taken of 3I/ATLAS on 2025 July 3. Two g-band exposures and two i-band exposures were taken with a total cumulative exposure time of 150 s, and three r-band exposures were taken with a total cumulative exposure time of 270 s. Additionally, a sidereal-tracked image containing 3I/ATLAS was taken to provide star detections to compare with the surface brightness profile of 3I/ATLAS. The g, r, and i images were taken over an airmass range of 1.79 to 2.05 during the 2025 July 3 observations. The P200 is located on Palomar Mountain, California.
@@
Name | g–r4 | r–i5 | g–i6 | Af7 | 8 | Refs. | |||
(au) | (deg) | (cm) | (kg/s) | ||||||
1I/’Oumuamua | 0.26 | 1.20 | 122.74 | 0.840.05 | 0.360.1 | 1.200.11 | 0 | 0 | Meech et al. (2017); Micheli et al. (2018) |
2I/Borisov | 2.01 | 3.36 | 44.05 | 0.630.05 | 0.200.02 | 0.830.05 | 14310 | 1 | Fitzsimmons et al. (2019); Jewitt & Luu (2019) |
3I/ATLAS | 1.35 | 6.08 | 175.11 | 0.430.02 | 0.160.02 | 0.590.03 | 2873 | 1 | This work. |

### v3

@@ Abstract
We present the characteristics of interstellar comet 3I/ATLAS, discovered on 2025 July 1 by the Asteroid Terrestrial-impact Last Alert System. The comet has eccentricity 6.08 and velocity at infinity 57 km/s, indicating an interstellar origin. We obtained B,V, R, I, g, r, i, and z photometry at Kottamia Astronomical Observatory, Palomar Observatory, and Apache Point Observatory on 2025 July 2, 3, and 6. We measured colour indices B–V=0.980.23, V–R=0.710.09, R–I=0.140.10, g–r=0.840.05, r–i=0.160.03, i–z=–0.020.07, g–i=1.000.05, and a spectral slope of 16.01.9 /100 nm. We estimate the comet’s dust cross-section within 10,000 km to be 184.64.6 km2, assuming a 0.1 albedo. 3I/ATLAS’s coma has FWHM2.2 arcsec and A(0∘)f=280.83.2 cm. We estimate 3I/ATLAS is ejecting µm to mm dust at 0.01-1 m/s, with a mass loss of 0.1 - 1.0 kg/s.
@@ 2 Observations
Initially designated as A11pl3Z, interstellar comet 3I/ATLAS was discovered at the ATLAS Rio Hurtado, Chile facility (Minor Planet Center (MPC) observing code W68). The initial discovery observations of 3I/ATLAS were made on 2025 July 1 05:15:11 in the ATLAS o-band filter as seen in Panels a-d of Figure 1 (Denneau et al., 2025). The comet was discovered at right ascension, declination of 18:07:27.68 and -18:41:40.2 and was detected in four o-band exposures taken between 2025 July 1 05:15:11and 2025 July 1 06:20:31. A list containing details of the ATLAS observations is presented in Section S.1 and Table S1 of the Supplemental Material.
@@
Images from all three facilities were reduced using biases and flats taken for each of the filters. Image stacks of the NGPS and ARCTIC data are shown in the mosaic images in Figs. 2 and 3. The multi-band data were calibrated using the Pan-STARRS1 catalog (Chambers et al., 2016) and the colour transformations from (Tonry et al., 2012) to convert Pan-STARR1 photometry magnitudes into SDSS photometry magnitudes. The photometry of 3I/ATLAS was measured using a 4″ radius circular aperture (10,000 km from the comet’s nucleus at its 3.36 - 3.45 au distance from the Earth during our KAO 1.88-m, P200, and ARCTIC 3.5-m observaitons). We used a 6.5-7.8″ radius sky background annulus to measure our photomery. The seeing at the KAO 1.88-m, the P200, and the ARC 3.5-m measured between 1.3-1.6 arcsec in images containing the comet. A list of the details of the KAO 1.88-m, P200, and ARC 3.5-m observations is presented in Section S.1 and Table S1 of the Supplemental Material.
@@ 3.1 Orbit
The astrometry of 3I/ATLAS from ATLAS taken in the discovery images on 2025 July 1 was reported to the MPC. Astrometry of the comet in 319 observations taken from 97 observational facilities around the world between 2025 June 14.25 UTC and 2025 July 4 UTC was subsequently reported to the MPC 111https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=3I, accessed on 2025 July 4.. The orbital solution from JPL HORIZONS, using observations taken between June 14, 2025, and July 4, 2025, is shown in Table S2. The location and orbit of 3I/ATLAS at the time of its discovery on 2025 July 1 is shown on Figure 4 based on values from JPL Horizons222https://ssd.jpl.nasa.gov/horizons/app.html#/.
@@ 3.2 Visible photometry
We computed the brightness of 3I/ATLAS in the KAO 1.88-m images of B=18.90 0.21, V=17.92 0.01, R=17.22 0.05, and I=17.08 0.09, in the NGPS images of g=18.720.02, r=17.860.01, and i=17.720.02, and g=18.580.06, r=17.770.02, i=17.590.04, and z=17.610.05 in the ARCTIC images. We used the NGPS r-band stack to search for evidence of an extended coma in 3I/ATLAS. The surface brightness profile extracted from the r-band images of 3I/ATLAS (blue line in Panel D of Figure 2) shows evidence of being extended, having 2-10 times the flux compared to the surface brightness profile of a comparison star (orange line in the same panel) outside of 1.3 arcsec of both profiles. The FWHM of the coma in the NGPS r-band stack is 2.2 arcsec compared to the 1.6 arcsec FWHM of background stars in the sidereal-tracked images. 3I/ATLAS ’s colours are B–V=0.980.23, V–R=0.710.09, R–I=0.140.10, g–r=0.840.05 mag, r–i=0.160.03 mag, i–z=–0.020.07 mag, and g–i=1.000.05 mag corresponding to a spectral slope of 16.01.9 /100 nm., somewhat more red than interstellar comet 2I/Borisov which had g–r=0.630.05, r–i=0.200.02, g–i=0.830.05, and a spectral slope of 9.71.8 /100 nm. The colors of 3I/ATLAS derived from our observations are similar to those reported by other photometric and spectroscopic studies (e.g., Opitom et al., 2025). We plot the g–r, r–i, and r–z colours of 3I/ATLAS in Figure 5.
@@
where is the half-width half maximum measured from our r-band stack of 1.1 arcsec, is the measured half-width half maximum of background stars of 0.8 arcsec, resulting in a of 0.7 arcsec, which translates to 1850 km at the 3.43 au distance of the comet. Plugging our calculated value of of 1850 km into Eq. 4 gives 1.1m s-1 suggesting that m-scale to mm-scale particles are ejected from the comet at 0.01-1 m/s. The dust mass loss, can be roughly estimated from , A(0∘)f, the dust albedo , the radius of the dust particles, , and the dust density, , with the following relation derived from (Fink & Rubin, 2012):
@@ 4 Discussion and conclusion
Interstellar comet 3I/ATLAS exhibits an active appearance and colours similar to some solar system comets, as seen in Figures 2 and 5. However, it seems to be significantly more red in its color compared to the other interstellar comet 2I (g–r=0.630.05, r–i=0.200.05, g–i=0.830.05, and r–z=–0.030.06), however, is less red compared to 1I/’Oumuamua (g–r=0.840.05, r–i=0.310.09, g–i=1.150.10, and r–z=0.410.10). However, this latter comparison may be inaccurate due to the fact that 3I/ATLAS is an active body, and 1I/’Oumuamua was only very weakly active as evidenced by the visible lack of a coma (Micheli et al., 2018). A complete table of the physical properties of the different interstellar objects is included in Table S3. Interstellar comet 3I/ATLAS appears to be more red than the majority of comets as seen in Fig. 4. The increased redness of 3I compared to 2I and other active comets may be due to the nature of its activity if it lacks the emission of gas molecules present in the g-band wavelength such as CN, C2, and C2 (Farnham et al., 2000). With our small sample size of three objects, it appears that the interstellar objects span a wide range in color. Additionally, 3I shows evidence of having a similar dust mass loss compared to 2I, but has a much higher dust mass loss rate than 1I, which was observed to show a negligible amount of dust mass loss.
@@ Acknowledgements
et al., 1998). The seeing in r-band was 1.6 arcsec measured by determining the full-width at half-maximum of background stars during the observations taken of 3I/ATLAS on 2025 July 3. Two g-band exposures and two i-band exposures were taken with a total cumulative exposure time of 150 s, and three r-band exposures were taken with a total cumulative exposure time of 270 s. The photometry from the g images was measured in individual images, and combined. Additionally, a sidereal-tracked image containing 3I/ATLAS was taken to provide star detections to compare with the surface brightness profile of 3I/ATLAS. The g, r, and i images were taken over an airmass range of 1.79 to 2.05 during the 2025 July 3 observations. The P200 is located on Palomar Mountain, California.
@@
name | q | e | i | g-r | r-i | r-z | g-i | Af | Refs. | |
(au) | (deg) | (cm) | (kg/s) | |||||||
1I/’Oumuamua | 0.26 | 1.20 | 122.74 | 0.840.05 | 0.360.1 | 0.360.10 | 1.200.11 | 0 | 0 | Meech et al. (2017) |
2I/Borisov | 2.01 | 3.36 | 44.05 | 0.630.05 | 0.200.02 | –0.030.11 | 0.830.05 | 14310 | 1 | Bolin et al. (2020) |
3I/ATLAS | 1.35 | 6.08 | 175.11 | 0.840.05 | 0.160.03 | 0.140.08 | 1.000.05 | 280.83.2 | 1 | This work. |